Circular No. 8536

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 L2 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery position

below) with the 0.5-m Uppsala Schmidt in the course of the Siding Spring Survey; 60-s exposures on June 2.6 UT showed the images to appear soft and probably cometary in reasonably good seeing. Exposures on June 3.4 showed the object to be nearly stellar with a 10′′ narrow tail in p.a. 90^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
June 2.55561  13^{h}22^{m}35.^{s}46  −47^{o}47′28′′.6  18.5

The following preliminary parabolic orbital elements are from MPEC 2005-L31, where an ephemeris also appears:

T = 2005 July 8.504 TT  ω = 292.^{o}454
Ω = 155.523 }2000.0
q = 3.19950 AU  i = 152.809

COMET C/2005 L3 (McNAUGHT) R. H. McNaught has reported yet another discovery of a comet (dis-

covery position below) with the 0.5-m Uppsala Schmidt in the course of the Siding Spring Survey; 60-s exposures on June 3.8 UT showed a condensation of diameter 10′′, while 20-s exposures on June 4.6 showed diameter ∼ 8′′ in worse seeing. Observations by A. C. Gilmore (Mt. John Observatory, 1.0-m f/7.7 reflector) on June 4.6 showed the object to be slightly more diffuse than the surrounding stars, although the object’s nature was unclear in poor seeing on June 5.6.

2005 UT  α_{2000}  δ_{2000}  Mag.
June 3.67556  21^{h}36^{m}01.^{s}71  −33^{o}05′00′′.3  17.7

The initial astrometric observations, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-L17.

T = 2008 Jan. 10.955 TT  ω = 50.^{o}099
Ω = 289.093 }2000.0
q = 5.33517 AU  i = 138.781
2005 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green